OpenStudy (anonymous):

Write 0.123 as a fraction. (123 has line over it)

OpenStudy (anonymous):

you want the quick answer or the slow one?

OpenStudy (anonymous):

It doesn't matter quick or slow as long as I can understand. :)

OpenStudy (anonymous):

quick way you have three digits that repeat \[\frac{123}{999}\]which reduces

OpenStudy (anonymous):

two digits repeating like \[12121212...=\frac{12}{99}\]

OpenStudy (anonymous):

one digit that repeats like \[.3333...=\frac{3}{9}\]

OpenStudy (anonymous):

understandable enuf, or you want the slow way?

OpenStudy (anonymous):

Understandable. Thank you.

OpenStudy (anonymous):

yw